OpenStudy (highschoolmom2010):

The measure of two complementary angles are in the ratio 1 : 5. What are the degree measures of the two angles? What is the solution of each proportion? A. 36° and 144° B. 15° and 75° C. 30° and 150° D. 18° and 72°

OpenStudy (highschoolmom2010):

i think it is C

OpenStudy (highschoolmom2010):

but i got it marked wrong could it be any of the other choices?

OpenStudy (highschoolmom2010):

ok... so we have \[36*5\neq144\] \[15*5=75\] ^ true \[30*5=150\] ^true \[18*5\neq72\]

HanAkoSolo (jamierox4ev3r):

it seems as if it would be B. because: \(\frac{1}{5}\)=\(\frac{15}{75}\)

HanAkoSolo (jamierox4ev3r):

this is the proportion that represents your answer.

OpenStudy (highschoolmom2010):

could it also be C \[\frac{ 1 }{ 5 }=\frac{ 30 }{ 150 }\]

OpenStudy (highschoolmom2010):

5*30=150=1 150=150

OpenStudy (radar):

50 and 130 are not complementary.

OpenStudy (highschoolmom2010):

ah ok ty @radar

OpenStudy (anonymous):

Let one angle be x Other angle = 5x Given, x + 5x = 90 6x = 90 x = 15 Other angle is 75